OpenStudy (anonymous):

For the following equation, tell which of the given ordered pairs are solutions. (Select all that apply.) 8x-7y=56 (0, -8) (8, 8) (0, 0) (8, 1) (7, 0)

OpenStudy (anonymous):

you just need to sub your answer choices into the equation and see if it is true. I will do the first two for you.... (1) (0,-8) x = 0 and y = -8 now sub 8x - 7y = 56 8(0) - 7(-8) = 56 56 = 56 (true..so this is a solution) (2) (8,8) x = 8 and y = 8 sub 8x - 7y = 56 8(8) - 7(8) = 56 64 - 56 = 56 8 = 56 (false, so this is not a solution) thats how you do it :) now you can finish

OpenStudy (anonymous):

thank you